OpenStudy (fearbigjjrob):

Can I get some quick help on Algerbra, I have the answer I believe but not sure.... question and answer in comments

OpenStudy (fearbigjjrob):

Solve 15=2^(x+2) how is s^2 translated if I change it to 2^(x+3) I believe it is either x+1=logV215 and for x+3; x+3=logV215 question 2 simplify into one logarithm: 2 log x + 3 log y I got 5 logV10xy

OpenStudy (fearbigjjrob):

@Nnesha

OpenStudy (fearbigjjrob):

@Jaynator495

OpenStudy (fearbigjjrob):

@griefer1010 ?

OpenStudy (fearbigjjrob):

@hartnn